### v2

# CodeUltraFeedback: An LLM-as-a-Judge Dataset for Aligning Large Language Models to Coding Preferences

Abstract.
Evaluating the alignment of large language models (LLMs) with user-defined coding preferences is a challenging endeavour that requires a deep assessment of LLMs’ outputs. Existing methods and benchmarks rely primarily on automated metrics and static analysis tools, which often fail to capture the nuances of user instructions and LLM outputs. To address this gap, we propose using the LLM-as-a-Judge methodology to evaluate the alignment of LLMs with coding preferences. Based on this approach, we present CodeUltraFeedback, a comprehensive dataset designed to facilitate the evaluation and improvement of LLM alignment. CodeUltraFeedback consists of 10,000 coding instructions, each annotated with four responses generated from a diverse pool of 14 LLMs. These responses are ranked based on five distinct coding preferences using GPT-3.5 as a judge, providing both numerical scores and detailed textual feedback. Our analysis of CodeUltraFeedback reveals that responses from GPT-3.5 and GPT-4 are generally preferred over those from open-weight LLMs, highlighting significant differences in alignment between closed and open-weight models. In turn, we explore the usage of CodeUltraFeedback as feedback data to fine-tune and align CodeLlama-7B-Instruct using supervised fine-tuning (SFT) and reinforcement learning from AI feedback (RLAIF) with direct preference optimization (DPO). The resulting aligned CodeLlama-7B-Instruct model outperforms larger LLMs in terms of alignment with coding preferences and shows improved functional correctness on the HumanEval+ benchmark compared to the original instruct model. Therefore, our contributions bridge the gap in preference tuning of LLMs for code and set the stage for further advancements in model alignment and RLAIF in automated software engineering.
Large language models, code generation, automated software engineering, reinforcement learning from AI feedback, direct preference optimization
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
The advent of recent large language models (LLMs) has ushered in a new era of LLMs with high coding capabilities (Chen and others., 2021; Roziere et al., 2023; Guo et al., 2024; Jiang et al., 2023; Touvron et al., 2023; Muennighoff et al., 2023), showcasing remarkable performances across a wide range of downstream tasks including code generation (Liu et al., 2023b; Athiwaratkun et al., 2022; Weyssow et al., 2023; Zheng et al., 2023b), code translation (Jiao et al., 2023; Pan et al., 2023), bug fixing (Silva et al., 2023; Muennighoff et al., 2023; Ye et al., 2021), and more. However, as LLMs continue to advance, a critical question emerges: How well do these capabilities align with developers’ expectations, particularly regarding non-functional requirements such as code readability, efficiency, and adherence to best practices?
Current methodologies for fine-tuning and evaluating LLMs primarily focus on core capabilities, e.g., translating, summarizing, or reviewing code (Lu et al., 2021; Puri et al., 2021; Zhu et al., 2022; Niu et al., 2023; Zhou et al., 2023; Sghaier and Sahraoui, 2024; Sghaier et al., 2023; Sghaier and Sahraoui, 2023), and functional correctness in diverse code generation scenarios (Hendrycks et al., 2021; Austin et al., 2021; Chen and others., 2021; Muennighoff et al., 2023; Liu et al., 2023b; Khan et al., 2023). While some research has started to address non-functional properties of LLM-generated code, such as code quality (Siddiq et al., 2023), runtime efficiency (Huang et al., 2024), and other aspects (Yetiştiren et al., 2023; Singhal et al., 2024), these efforts often fall short in several key areas. First, existing benchmarks typically refine established metrics without considering more open-ended coding problems that involve complex user instructions. Second, the assessment of LLM outputs often relies on automated metrics and external tools based on inflexible standards and patterns. This approach fails to capture the nuanced complexity of users’ instructions and the subtleties in LLMs’ responses. Third, there is a notable absence of large-scale datasets specifically designed for tuning and aligning LLMs to non-functional requirements.
The landscape of existing datasets and benchmarks in automated software engineering reveals a critical gap: the lack of methods for both tuning LLMs and accurately measuring their alignment with non-functional requirements. Furthermore, current evaluation methods fall short of capturing the nuances in natural and programming languages, underscoring the need for a more sophisticated, language-centric approach to assessment. To address these shortcomings, we must develop comprehensive datasets designed to improve LLM alignment with non-functional requirements and design evaluation methods capable of evaluating LLM alignment with greater precision. By closing these gaps, we can better bridge the divide between developers’ expectations and the output of LLMs, ensuring that the generated code not only functions correctly but also adheres to best practices, maintains readability, and operates efficiently.
In this paper, we introduce two novel contributions: CodeUltraFeedback and CodeUltraFeedback-Bench. CodeUltraFeedback is a preference dataset comprising 10,000 complex instructions with 40,000 meticulously curated LLMs responses aligned with five non-functional requirements (or coding preferences [^1] In the rest of this paper, we use the term “coding preferences” instead of “non-functional requirements” to more precisely capture the process of instructing LLMs with explicit preferences, although both terms bear similar meanings.): instruction following, code explanation, code complexity and efficiency, code readability, and coding style. The objective of CodeUltraFeedback is to serve as a dataset for preference tuning of LLMs, leveraging recent advancements in LLM alignment, including UltraFeedback (Cui et al., 2023), reinforcement learning from AI feedback (RLAIF) (Bai et al., 2022; Lee et al., 2023; Ouyang et al., 2022) and LLM-as-a-Judge (Zheng et al., 2023a; Li et al., 2023a), predicated upon the advanced judging capabilities of LLMs like GPT-3.5 or GPT-4. We adopt an approach auxiliary to UltraFeedback for building our dataset, starting by tagging each instruction with a coding preference. Then, we generate responses to the instructions using four LLMs randomly selected from a pool of 14 LLMs to achieve diversity and consider various writing styles. Finally, we use LLM-as-a-Judge with GPT-3.5 to rate the alignment of the LLM responses with the instruction’s coding preference, resulting in annotations comprising a numerical rating and a textual rationale for the rating. The resulting dataset can then be leveraged to align an LLM to generate high-quality responses given a coding preference by tuning the LLM through RLAIF using techniques like direct preference optimization (DPO) (Rafailov et al., 2023).
Subsequently, we construct CodeUltraFeedback-Bench by selecting a subset of 500 instructions from CodeUltraFeedback. CodeUltraFeedback-Bench serves a different purpose than CodeUltraFeedback and aims to comprehensively assess and compare the alignment of LLMs with coding preferences. We design a rigorous single-answer grading scheme using LLM-as-a-Judge that allows the judge LLM to evaluate LLM-generated code consistently and objectively. This approach offers a more nuanced evaluation strategy than previous methods that rely on automated metrics and external tools by leveraging the advanced reasoning capabilities of LLMs. Their ability to discern nuances in language allows for a more refined and context-sensitive evaluation of how well LLM-generated code aligns with coding preferences in a language-centric fashion.
An initial data analysis of CodeUltraFeedback’s annotations reveals a lack of alignment of 12 LLMs, including strong LLMs such as WizardCoder-33B (Luo et al., 2023) and DeepSeek-Coder-Instruct-33B (Guo et al., 2024), while GPT-3.5 and GPT-4 demonstrate strong alignment. In turn, we explore preference tuning of a small model, CodeLlama-7B-Instruct (Roziere et al., 2023) using CodeUltraFeedback with supervised fine-tuning (SFT) and RLAIF with DPO (Rafailov et al., 2023; Tunstall et al., 2023). This method selects high and low-rated responses to align CodeLlama-7B-Instruct with coding preferences using DPO, encouraging the model to favor highly rated responses. We demonstrate that tuning CodeLlama-7B-Instruct with SFT and DPO substantially improves alignment across all coding preferences, surpassing all larger LLMs, including CodeLlama-34B-Instruct and WizardCoder-33B when using GPT-3.5-Turbo as a judge. Further, we explore LLM alignment on CodeUltraFeedback-Bench using GPT-4-Turbo as a judge and highlight more mitigated effects of SFT and DPO, albeit CodeLlama-7B-Instruct still outperforms CodeLlama-13B-Instruct and CodeLlama-34B-instruct in this setup. We also show that preference tuning does not hinder CodeLlama-7B-Instruct’s ability to generate functionally correct code, substantially improving Pass@$k$ on HumanEval (Chen and others., 2021) and HumanEval+ (Liu et al., 2023b). Finally, we implement QLoRA (Dettmers et al., 2023) for efficient fine-tuning, showing that SFT and DPO can effectively operate on a single RTX A5000 GPU (24GB). Our work establishes strong evidence of the benefits of SFT and DPO in improving LLM alignment with coding preferences, demonstrating advancements in both preference alignment and functional correctness. This paves the way for future improvements in LLM training and evaluation, promising closer alignment with developers’ nuanced expectations.
To summarize, our main contributions are the following:
  • We release CodeUltraFeedback, a preference dataset of 10,000 complex instructions and 40,000 responses generated using 14 diverse LLMs for aligning LLMs to coding preferences in a code generation scenario.
  • We contribute CodeUltraFeedback-Bench, a benchmark to evaluate and compare LLM alignment using LLM-as-a-Judge over five coding preferences: instruction following, code explanation, code complexity and efficiency, code readability, and coding style.
  • We validate the utility of CodeUltraFeedback in facilitating LLM alignment of CodeLlama-7B-Instruct using SFT and DPO.
Our models, dataset, benchmark, and prompt templates are available at https://github.com/martin-wey/CodeUltraFeedback.
2. CodeUltraFeedback
In this section, we introduce our new dataset, CodeUltraFeedback. Fig. 1 depicts our methodology for building CodeUltraFeedback, which incorporates various components and ideas that we describe in the subsequent subsections.
2.1. Coding Preferences and Principles.
We start by identifying five coding preferences to guide the creation of our dataset and which are essential to evaluate the broader capabilities of LLMs: (1) Instruction-Following is about the strict adherence of the LLM to the instructions provided by users. This preference is foundational for ensuring that LLMs truly follow the user intent and thus provide personalized responses to instructions. (2) Code Explanation emphasizes generating code with detailed natural language explanations. It underscores the importance of an LLM in providing a solution with explanations that can serve as a bridge between potentially complex code and users while improving trustworthiness. (3) Code Complexity and Efficiency preference underlines the LLM capability to generate code optimized for performance in terms of speed and resource utilization. It is another crucial aspect requiring the LLM to balance the speed and resource usage of the solution carefully. (4) Code Readability serves a distinct purpose to the code explanation preference by emphasizing the clarity and understandability of the code itself through its structure, style, and the presence of meaningful documentation and in-line comments. (5) Coding Style preference focuses on the importance of writing code that not only meets syntactical correctness but also aligns with the idiomatic practices and stylistic norms of the programming language.
Altogether, these five preferences complement the functional correctness requirement of LLM-generated code. Assessing how well LLMs align with these preferences amounts to evaluating how the generated solutions are optimized for human comprehension, maintainability, performance, and collaborative work.
Figure 1. Overview of CodeUltraFeedback dataset construction procedure. (i) Coding Preferences and Principles: we define five coding preferences: Instruction-following, Code Explanation, Code Complexity and Efficiency, Code Readability, and Coding Style, and 10 principles for each preference to guide LLMs’ generative process. (ii) Magicoder Evol-Instruct Dataset: CodeUltraFeedback is based on a 10k-instruction subset of Magicoder Evol-Instruct dataset (Wei et al., 2023; Luo et al., 2023), where each instruction is tagged with a coding preference. (iii) LLMs Selection: for each instruction, four LLMs are randomly sampled from a diverse pool of 14 LLMs. (iv) Principle-Driven Code Generation: one principle is randomly selected per LLM to guide and align the code generation process with the assigned coding preference. (v) LLM-as-a-Judge Annotation: LLM-as-a-Judge with GPT-3.5-Turbo is used to judge LLMs’ responses according to coding preferences evaluation criteria (see Table 1).
2.2. Initial Dataset.
To build CodeUltraFeedback, we rely on Magicoder Evol-Instruct dataset (Wei et al., 2023), an Evol-Instruct (Xu et al., 2023; Luo et al., 2023) version of CodeAlpaca dataset comprising coding problem instructions. We select a random subset of Magicoder Evol-Instruct comprising 10,000 samples serving as an initial dataset for CodeUltraFeedback. We assign 2,000 samples per coding preference, ensuring a balanced representation of each preference in the dataset. The reason for subsampling the dataset to 10,000 samples is to lower costs related to OpenAI API needed to generate some of the responses and the annotations.
2.3. LLMs Selection.
In order to generate highly diverse solutions to the instructions, we select a pool of 14 LLMs spanning eight model families. These include GPT-4-Turbo and GPT-3.5-Turbo as closed-source LLMs. We include Llama-2-13B-Chat and Llama-2-70B-Chat (Touvron et al., 2023), specialized for chat use cases. We choose CodeLlama-7/13/34B-Instruct models (Roziere et al., 2023), instruction-tuned for code generation. We also include WizardLM-15/33B models (Xu et al., 2023) and Mistral-7B-Instruct (Jiang et al., 2023), which are instruction-tuned LLMs that can follow complex instructions. Finally, we include WizardCoder-15/33B (Luo et al., 2023) and DeepSeek-Coder-6.7/33B-Instruct (Guo et al., 2024), recent instruction-tuned LLMs that demonstrate state-of-the-art performance on code generation benchmarks such as HumanEval (Chen and others., 2021) and HumanEval+ (Liu et al., 2023b). From this pool, we randomly select four LLMs for each instruction to generate responses, allowing for the representation of a variety of coding approaches and styles.
2.4. Principle-Driven Code Generation.
In this step, we leverage principle-driven generation and generate 10 principles per coding preference using ChatGPT following prior work practices (Sun et al., 2023; Cui et al., 2023; Mukherjee et al., 2023). The rationale for establishing 10 principles per preference is to diversify the dataset. All principles for each coding preference can be found in Tables 1620.
During code generation, one principle, corresponding to the coding preference of the sample, is randomly chosen per LLM and appended to the instruction in an input prompt to guide the generation. This strategy helps the LLMs generate responses that potentially align with the requirements of the coding preferences. The output of this step is the four LLMs’ responses to the principle-augmented instructions.
2.5. LLM-as-a-Judge Annotation.
Table 1. Code readability assessment template.
Evaluate the readability of code segments. Assess how comments and documentation contribute to understanding the code’s logic, purpose, and operation. Evaluation Criteria:
   Clarity: How clear and understandable are the code and its accompanying comments/documentation?
   Conciseness: Are the comments and documentation succinct yet informative?
   Relevance: Do the comments and documentation directly contribute to explaining the code’s logic, objectives, and functionality?
   Comprehensibility: Can users of varying technical backgrounds easily grasp the code’s purpose and how it works?
Scoring: Rate outputs on a scale of 1 to 5:
  1. Poor Readability: The code is hard to follow, with little to no helpful comments/documentation.
  2. Basic Readability: The code has minimal comments/documentation, offering limited clarity or insight.
  3. Good Readability: The code is reasonably clear with comments/documentation that aid understanding, though some areas could be improved.
  4. Very Good Readability: The code and comments/documentation are clear and concise, making the code’s logic and purpose easily understandable.
  5. Excellent Readability: The code exemplifies outstanding readability, with clear, concise, and comprehensive comments/documentation that make it accessible to all users.
We apply the LLM-as-a-Judge methodology (Zheng et al., 2023a) to annotate responses using GPT-3.5-Turbo. Drawing from practices established in previous work (Cui et al., 2023), we implement a set of guidelines ensuring GPT-3.5-Turbo delivers both consistent and objective feedback. Each annotation consists of (1) a score on a 1–5 scale offering a comprehensive assessment of the response’s alignment, and (2) a rationale for the assigned score, highlighting the response’s strengths and identifying any potential shortcomings. Tasking GPT-3.5-Turbo to provide a rationale for the scores encourages a chain-of-thought (Wei et al., 2022) process during the generation of annotations, elucidating the reasons behind the scores and meaningful feedback. In addition, this process mitigates the risk of variability and subjectivity that might arise from purely quantitative assessments. Consequently, this reflective process guarantees that the scoring is not just a surface-level judgment but can be backed by thoughtful consideration of some well-defined coding preferences evaluation criteria.
Table 1 illustrates the assessment template and evaluation criteria for code readability evaluation. It aims to provide precise and comprehensive criteria to the LLM judge for response assessment. We design analogous assessment templates, each tailored with specific evaluation criteria, for evaluating the other coding preferences (see Tables 69). The annotation procedure entails constructing a prompt combining the coding preference assessment template and the four LLMs responses. GPT-3.5-Turbo is then tasked to generate the annotations for all four responses simultaneously, enhancing the consistency and reliability of the evaluation process over individual response assessments. Table 2 presents an example of the output of the annotation process using GPT-3.5-Turbo, with the LLMs responses and some rationales omitted for brevity. This example showcases GPT-3.5-Turbo’s capability to provide comprehensive justifications for each rating, demonstrating its effective evaluation capabilities. Furthermore, Table 2 illustrates an example of relatively complex instructions included in CodeUltraFeedback, showcasing the complexity of the task.
Table 2. Excerpt of ratings and rationales generated by GPT-3.5-Turbo using LLM-as-a-Judge prompting. GPT-3.5-Turbo generates clear rationales for each rating demonstrating its strong judging capabilities.
Instruction: Devising a Python-based algorithm that utilizes the technique of ternary search, could you draw this up while also illustrating the functionality of each critical stage in the searching process?
Preference: Readability
GPT-4: [Response]
Rating: 5
Rationale: The code is accompanied by clear and concise comments that comprehensively explain each critical stage of the ternary search algorithm. The comments provide a thorough understanding of the algorithm’s logic, purpose, and functionality, making it easily accessible to users of varying technical backgrounds. Additionally, the comments include a detailed explanation of the algorithm’s relevance and efficiency compared to binary search, enhancing the overall comprehensibility of the code.
DeepSeek-Coder-6.7B-Instruct: [Response]
Rating: 4
Rationale: [Rationale]
Mistral-7B-Instruct: [Response]
Rating: 3
Rationale: [Rationale]
CodeLlama-13B-Instruct: [Response]
Rating: 2
Rationale: The code is clear, but the comments are minimal and lack sufficient detail to fully explain the ternary search algorithm and its critical stages. While the comments briefly describe the function’s purpose, they do not provide a comprehensive explanation of the algorithm’s logic and functionality. Additionally, the comments do not cover the relevance of ternary search in algorithmic strategies, making it less accessible to users with varying technical backgrounds.
3. CodeUltraFeedback: Dataset Analysis
Table 3. Average scores across coding preferences for each LLM. The statistical significance of differences in ratings between each LLM and GPT-3.5-Turbo was tested using pairwise Welch’s $t$-tests. Significance levels are denoted as follows: * for $p<0.05$, $\dagger$ for $p<0.01$, and $\ddagger$ for $p<0.001$. The results from almost all pairwise $t$-tests indicate highly significant differences ($p<0.001$).
Model Instruction Following Code Explanation Code Complexity & Efficiency Code Readability Coding Style Average
GPT-4-Turbo $\mathbf{3.79}^{\ddagger}$ $\mathbf{4.04}^{\ddagger}$ $\mathbf{3.91}^{\dagger}$ $\mathbf{4.14}^{\ddagger}$ $\mathbf{4.03}^{\ddagger}$ 3.98
GPT-3.5-Turbo 3.56 3.76 3.76 3.71 3.66 3.69
WizardCoder-33B $3.29^{\ddagger}$ $3.31^{\ddagger}$ $3.43^{\ddagger}$ $3.44^{\ddagger}$ $3.49^{\dagger}$ 3.39
DeepSeek-Coder-33B-Instruct $3.31^{\ddagger}$ $3.30^{\ddagger}$ $3.32^{\ddagger}$ $3.46^{\ddagger}$ $3.42^{\ddagger}$ 3.36
DeepSeek-Coder-6.7B-Instruct $3.23^{\ddagger}$ $3.29^{\ddagger}$ $3.32^{\ddagger}$ $3.45^{\ddagger}$ $3.48^{\ddagger}$ 3.36
Mistral-7B-Instruct $3.20^{\ddagger}$ $3.27^{\ddagger}$ $3.28^{\ddagger}$ $3.42^{\ddagger}$ $3.40^{\ddagger}$ 3.31
CodeLlama-34B-Instruct $3.11^{\ddagger}$ $3.20^{\ddagger}$ $3.21^{\ddagger}$ $3.35^{\ddagger}$ $3.23^{\ddagger}$ 3.22
Llama-2-70B-Chat $3.11^{\ddagger}$ $3.22^{\ddagger}$ $3.14^{\ddagger}$ $3.38^{\ddagger}$ $3.22^{\ddagger}$ 3.21
WizardCoder-15B $3.11^{\ddagger}$ $3.03^{\ddagger}$ $3.12^{\ddagger}$ $3.27^{\ddagger}$ $3.16^{\ddagger}$ 3.14
CodeLlama-13B-Instruct $3.05^{\ddagger}$ $2.99^{\ddagger}$ $3.15^{\ddagger}$ $3.18^{\ddagger}$ $3.25^{\ddagger}$ 3.12
CodeLlama-7B-Instruct $2.91^{\ddagger}$ $3.11^{\ddagger}$ $3.01^{\ddagger}$ $3.18^{\ddagger}$ $3.13^{\ddagger}$ 3.07
WizardLM-33B $3.07^{\ddagger}$ $2.98^{\ddagger}$ $2.91^{\ddagger}$ $3.11^{\ddagger}$ $3.10^{\ddagger}$ 3.03
Llama-2-13B-Chat $2.88^{\ddagger}$ $2.97^{\ddagger}$ $2.91^{\ddagger}$ $3.18^{\ddagger}$ $2.99^{\ddagger}$ 2.98
WizardLM-7B $2.63^{\ddagger}$ $2.61^{\ddagger}$ $2.51^{\ddagger}$ $2.69^{\ddagger}$ $2.64^{\ddagger}$ 2.62
In this section, we provide a detailed analysis of CodeUltraFeedback. Specifically, we explore the inherent capabilities of LLMs to align with coding preferences. Furthermore, we showcase the diversity of the dataset in terms of LLM responses and scores.
3.1. LLMs Scores Exploration
We start by exploring the ratings generated by GPT-3.5-Turbo in the LLM-as-a-Judge annotation phase. In Table 3, we present a detailed analysis of the average scores across coding preferences for each LLM, offering initial insight into their baseline capabilities.
At first glance, the scores might suggest marginal discrepancies given the 1–5 scoring range. However, these variations are substantial and carry statistical significance. To quantify this significance, we performed pairwise Welch’s $t$-tests, comparing the score distribution for each LLM against GPT-3.5-Turbo across each coding preference. The results reveal highly significant statistical differences, with $p<0.001$, for most comparisons and validate the discernible performance gaps between all models and GPT-3.5-Turbo. Moreover, the results also demonstrate a notably superior performance of GPT-4-Turbo relative to GPT-3.5-Turbo.
Overall, GPT-3.5-Turbo and GPT-4-Turbo score the highest across all preferences and on average. This outcome is expected considering both models have undergone extensive instruction and reinforcement learning from human feedback (RLHF) tunings to align more closely with human preferences (Ouyang et al., 2022). The scores indicate that highly capable LLMs such as WizardCoder-33B and DeepSeek-Coder-33B-Instruct underperform compared to GPT-3.5-Turbo. This underperformance is likely due to a lack of alignment of the LLMs, and tuning these models using alignment techniques could improve their performance.
3.2. LLMs Responses Analysis
Figure 2. Percentage of best and worst responses per LLM.
Figure 3. Distribution of best and worst responses length.
To gain more insights into the diversity of CodeUltraFeedback, we analyze the best and worst responses for each sample in terms of their assigned scores.
Fig. 2 depicts the percentage of best and worst responses per LLM in the dataset. Although GPT-4-Turbo and GPT-3.5-Turbo consistently score the highest across all coding preferences on average, this plot highlights that another LLM response is favoured for 73% of the samples. Moreover, even LLMs with lower average scores produce the highest-rated responses for at least 3.1% of the samples. This property of our dataset is valuable for downstream applications. For instance, it enables the alignment of LLMs using feedback data that encompasses a diverse range of models and mitigates the risk of the aligned LLM mimicking the behaviour of a single model.
Next, we analyze the distribution of best and worst response lengths in Fig. 3. Both distributions show similar patterns, though the best responses tend to be slightly longer. Importantly, both distributions encompass a broad spectrum of response lengths. This characteristic of CodeUltraFeedback is also valuable for model alignment, as it helps prevent the aligned LLM from developing a bias towards excessive verbosity.
3.3. LLMs vs GPT-3.5-Turbo
To gauge the relative performance of the LLMs against GPT-3.5-Turbo more in-depth, we analyze their performances through a win-tie-lose plot depicted in Fig. 4. The idea is to pit each LLM (left-hand side of the figure, e.g., Mistral-7B-Instruct) against GPT-3.5-Turbo in a comparative match-up, where the model achieving the highest rating wins. The plot illustrates the percentage of ratings of an LLM that are higher/lower compared to those of GPT-3.5-Turbo, all preferences combined (a tie stands for ratings having identical values). This figure further demonstrates significant gaps between LLMs and GPT-3.5-Turbo, underscoring a consistent preference for GPT-3.5-Turbo and GPT-4-Turbo’s responses over other LLMs. For instance, even Mistral-7B-Instruct, the second-best LLM, achieves a win rate of merely 29.8%. Additionally, the plot further validates the utilization of GPT-3.5-Turbo as a judge, showcasing its capability to discern between responses of differing quality. For example, it is proficient in recognizing higher-quality responses, as evidenced by its comparison with GPT-4-Turbo, which boasts a win rate of 51.6%.
In conclusion, these initial observations highlight a significant research opportunity and the necessity to delve deeper into aligning LLMs with coding preferences to make them more competitive with models like GPT-3.5 and GPT-4.
Figure 4. Win-tie-lose ratios of LLMs against GPT-3.5-Turbo. An LLM wins/loses if it gets a greater/lower score than GPT-3.5-Turbo. A tie is when the LLM and GPT-3.5-Turbo get identical scores.
4. Aligning LLMs to Coding Preferences
In this section, we present our experimental setup to improve LLM alignment with coding preferences using CodeUltraFeedback as preference data, which features the utilization of SFT and DPO (Liu et al., 2023b; Tunstall et al., 2023). The purpose is twofold: validate the utility of CodeUltraFeedback for LLM alignment to coding preferences and show that a small LLM tuned using SFT and DPO can achieve greater alignment performance.
4.1. Supervised Fine-Tuning (SFT)
As demonstrated in Zephyr’s paper (Tunstall et al., 2023), an instruction-tuning phase is required prior to tuning an LLM using DPO as it facilitates tuning. This phase is achieved through supervised fine-tuning (SFT) using a dataset $\mathcal{D}=\{(x_{1},y_{1}),...,(x_{n},y_{n})\}$ of instruction-response pairs. Recent work leverages the Self-Instruct (Wang et al., 2022a) and Evol-Instruct (Xu et al., 2023; Luo et al., 2023) frameworks to compose such a dataset by generating instructions and responses using a highly capable model such as GPT-3.5/4 (Ouyang et al., 2022; Taori et al., 2023; Wei et al., 2023). In this context, SFT is a form of knowledge distillation (Xu et al., 2024) that leverages a supervised signal from responses generated by a teacher LLM to tune a student LLM.
Formally, for each instruction-response pair $(x_{i},y_{i})\in\mathcal{D}$, the learning objective is to minimize the cross-entropy loss $\mathcal{L}_{SFT}$, defined as:
$\mathcal{L}_{SFT}=-\frac{1}{n}\sum_{i=1}^{n}\log P(y_{i}|x_{i},\theta),$
where $P(y_{i}|x_{i},\theta)$ represents the probability of generating the response $y_{i}$ given the instruction $x_{i}$, parameterized by $\theta$, the LLM parameters.
4.2. Direct Preference Optimization (DPO)
DPO is an efficient ranking optimization method to align LLMs to preference data, where a response $y_{w}$ is preferred over a rejected response $y_{l}$. Unlike traditional reinforcement learning methods like RLHF, which require training a separate reward model, DPO enhances stability and performance by directly adjusting the LLM’s policy towards higher-ranked responses (i.e., $y_{w}$). Given a dataset $\mathcal{P}$ of triplets $(x,y_{w},y_{l})$, a reference LLM policy $\pi_{ref}$, and an LLM policy $\pi_{\theta}$, the objective is to maximize the following expectation:
$\mathbb{E}_{\begin{subarray}{c}(x,y_{w},y_{l})\sim\mathcal{P}\end{subarray}}\log\sigma\left(\beta\log\left(\frac{\pi_{\theta}(y_{w}|x)}{\pi_{ref}(y_{w}|x)}-\frac{\pi_{\theta}(y_{l}|x)}{\pi_{ref}(y_{l}|x)}\right)\right).$
The expression $\left(\frac{\pi_{\theta}(y_{w}|x)}{\pi_{ref}(y_{w}|x)}-\frac{\pi_{\theta}(y_{l}|x)}{\pi_{ref}(y_{l}|x)}\right)$ calculates the difference in probabilities that the model assigns to the preferred response $y_{w}$ and the response $y_{l}$, relative to the reference policy $\pi_{ref}$. Intuitively, the difference indicates how much the preference alignment of $\pi_{\theta}$ has improved for a given triplet $(x,y_{w},y_{l})$ after applying DPO compared to the reference policy $\pi_{ref}$. $\beta$ is a hyperparameter that adjusts the sensitivity to the reference policy $\pi_{ref}$. In our experiments, $\pi_{ref}$ is a SFT-tuned LLM.
4.3. Experimental Details
Datasets. We build CodeUltraFeedback-Bench, a new benchmark for assessing LLM alignment to coding preferences. CodeUltraFeedback-Bench consists of 500 randomly selected samples from CodeUltraFeedback, with a balanced representation of each preference. The benchmark aims to provide a rigorous and comprehensive framework to evaluate and compare LLM alignment to coding preferences, allowing researchers to evaluate the impact of new alignment methods.
For SFT, we use Magicoder Evol-Instruct (Wei et al., 2023), which consists of instruction-response pairs generated by GPT-4. We filter out the 10,000 samples used to build CodeUltraFeedback to mitigate data leakage between the SFT and DPO phases and avoid potential overfitting. The process results in 100,772 samples (95% for training, 5% for model evaluation).
We use the remaining 9,500 samples from CodeUltraFeedback for DPO. From each sample, we extract binary preferences ($y_{w}$, $y_{l}$), denoting the preferred and rejected response. These preferences are selected based on the highest and lowest ratings assigned by GPT-3.5-Turbo during the LLM-as-a-Judge annotation phase (see Section 2.5). In instances where multiple responses have the highest/lowest ratings, we select one randomly to obtain $y_{w}$ / $y_{l}$.
LLMs. We conduct our experiments on CodeLlama-7B-Instruct as it comprises 7B parameters, allowing for tuning on a modest computing infrastructure. Furthermore, the model lags behind more proficient LLMs in CodeUltraFeedback’s scores (see Table 3), which makes it an ideal candidate for demonstrating the potential impact of SFT and DPO on improving LLM alignment. For comparison, we include the following LLMs: CodeLlama-13/34B-Instruct, DeepSeek-Coder-6.7B-Instruct, WizardCoder-33B, GPT-3.5-Turbo, and GPT-4-Turbo.
Grading Procedure. CodeUltraFeedback-Bench seeks to systematically compare LLM alignment, whereas CodeUltraFeedback’s main purpose is to serve as a training dataset for preference tuning. Therefore, we designed a new grading procedure adapted for evaluating each LLM individually using LLM-as-a-Judge. We implement a reference-guided single-answer grading system that has proven efficient in alignment benchmarks such as MT-Bench (Zheng et al., 2023a). Each LLM’s response is individually evaluated against a reference response, providing a consistent basis for comparison and ensuring objective judgments across LLMs. Additionally, we instruct the judge LLM to provide a rationale alongside the rating, enhancing the consistency of the evaluations. Ratings are assigned on a scale from 1 to 10, facilitating nuanced assessments of LLM alignment. We generate reference responses using GPT-3.5-Turbo and GPT-4-Turbo and use both models as judges.
Training Details. We use HuggingFace TRL (Leandro et al., 2020) to implement SFT and DPO with QLoRA (Dettmers et al., 2023) for efficient fine-tuning. We use a cosine learning rate scheduler with a learning rate of 2e-04 for SFT and 5e-05 for DPO, and 10% warmup steps. We fine-tune CodeLlama-7B-Instruct for 3 and 5 epochs for SFT and DPO, respectively. All experiments were conducted on a single NVIDIA RTX A5000 GPU (24GB). We provide extensive training details in our replication package.
5. Results
Figure 5. Average alignment scores for LLMs across coding preferences on CodeUltraFeedback-Bench, evaluated using GPT-3.5-Turbo as a judge with reference-guided single-answer grading.
SFT and DPO Improve Alignment to Coding Preferences. In Fig. 5, we report the average alignment scores of the LLMs on the five coding preferences using GPT-3.5-Turbo as a judge with its own responses to the instructions as references. We do not include GPT-3.5-Turbo in the results as references and responses would be identical, leading to biased and perfect alignment scores.
GPT-4-Turbo stands out by surpassing all other LLMs across every coding preference by a considerable margin. Interestingly, all LLMs demonstrate lower scores on the instruction following preference. This trend suggests a relative challenge in achieving high performance in this area, possibly due to the inherently less precise nature of the preference and potentially vague instructions. Conversely, LLMs tend to achieve higher scores on code explanation and readability preferences. We hypothesize this trend can be attributed to the fact that all LLMs have undergone instruction tuning, which can encourage LLMs to output reasoning and explanations alongside generated code. We observe that scores substantially increase when tuning CodeLlama-7B-Instruct with SFT+DPO, highlighting the effectiveness of these methods in enhancing model alignment across all coding preferences. For instance, the model yields a $7.89$ score on code complexity and efficiency preference, substantially surpassing CodeLlama-7B-Instruct ($6.16$) and larger models, including Wizard-33B ($6.93$) and CodeLlama-34B-Instruct ($6.42$).
Tuning CodeLlama-7B-Instruct with SFT shows an improvement in alignment scores compared to the base model, which are further elevated with the application of DPO independently or in conjunction with SFT. To quantify these observations, we conducted pairwise $t$-tests (not reported for brevety) comparing the baseline model, against its variants tuned with SFT, DPO, and SFT+DPO. Although SFT alone improves performance, we found no significant statistical difference with the baseline across preferences, except for code readability. However, we found highly significant statistical differences ($p<0.001$) for DPO and SFT+DPO for all preferences with the exception of instruction following. After DPO and SFT+DPO tunings, CodeLlama-7B-Instruct achieves superior alignment compared to substantially larger models such as CodeLlama-34B-Instruct and WizardCoder-33B across all coding preferences.
DPO Enhances Competitive Edge Against GPT-4-Turbo. Fig. 6 illustrates the win-tie-lose ratios of the LLMs against GPT-4-Turbo for all preferences combined.
Figure 6. Win-tie-lose ratios of LLMs against GPT-4-Turbo. DPO and SFT+DPO substantially increase the percentage of won and tie matches.
Non-aligned models achieve low win rates, with CodeLlama-7B-Instruct and WizardCoder-33B peaking at a 10.8% win rate. GPT-4-Turbo, on the other hand, consistently outperforms most LLMs with win rates exceeding 60%, underscoring its superiority. Tuning CodeLlama-7B-Instruct with DPO and SFT+DPO results in increases in win rates by 4.8% and 6.2%, respectively, and raises tie scores by 15.4% and 16.4%, respectively. These improvements underscore the effectiveness of alignment techniques in rendering LLM responses more competitive. Nonetheless, the remaining discernible gaps with GPT-4-Turbo suggest that there is still substantial room for improvement.
Different Judges and References Impact Alignment Scores. We investigate the effect of judges and references on the LLMs alignment scores in Table 4.
Using GPT-3.5-Turbo as a judge, alignment scores are compared against its own (left values) and GPT-4-Turbo’s responses (right values), revealing a decrease against the latter due to GPT-4-Turbo’s higher response standards. Additionally, CodeLlama-7B-Instruct+SFT+DPO achieves an alignment score of 7.08, nearing GPT-3.5-Turbo’s average of 7.18.
Interestingly, divergent trends emerge under GPT-4-Turbo’s judgment, with all CodeLlama models scoring lower, while other LLMs see score increases. For instance, CodeLlama-13B-Instruct’s score falls from 5.58 to 4.83, while WizardCoder-33B’s increases from 6.26 to 6.75. Despite this, SFT, DPO, and SFT+DPO alignment techniques boost CodeLlama-7B-Instruct’s performance, still exceeding the 34B variant of CodeLlama, albeit with varying impact with DPO.
These observations hint at GPT-4-Turbo’s potential preference for certain response styles and variability in judgment compared to GPT-3.5-Turbo, underscoring the need for further investigation to fully understand the implications of judge selection on LLM alignment scores.
SFT and DPO Improve Functional Correctness. While our findings highlight the positive impact of SFT and DPO on aligning LLMs with coding preferences, ensuring the functional correctness of the generated code remains a pivotal concern. In Table 5, we report CodeLlama-7B-Instruct’s Pass@$k$ on HumanEval (Chen and others., 2021) and HumanEval+ (Liu et al., 2023b).
Firstly, SFT substantially enhances Pass@$k$ on both benchmarks, with Pass@$1$ rising from 37.9 to 51.2 and 33.2 to 45.6 for HumanEval and HumanEval+, respectively. This result aligns with Magicoder (Wei et al., 2023) and WizardCoder (Luo et al., 2023) findings, showing that knowledge distillation through SFT substantially enhances the LLM’s effectiveness on these benchmarks. Secondly, while DPO and SFT+DPO variants exhibit a slight reduction in Pass@$k$ relative to SFT alone, they still maintain substantially higher Pass@$k$ compared to the baseline.
These results show that LLM alignment through SFT and DPO also improves functional correctness, mostly due to the SFT phase, further demonstrating the usefulness of aligning LLMs. The lesser improvements in Pass@$k$ with DPO can be attributed to DPO’s learning objective, focusing on favouring preferred responses with respect to coding preferences rather than based on their functional correctness. Therefore, we underscore the need for more investigations into the design of learning objectives and methods that can prioritize both functional correctness and alignment to coding preferences.
Table 4. Average alignment scores of LLMs on CodeUltraFeedback-Bench. G-3.5 and G-4 refer to utilizing GPT-3.5-Turbo and GPT-4-Turbo responses as references, respectively.
Judge
GPT-3.5-Turbo GPT-4-Turbo
                          Reference G-3.5 G-4 G-4
CodeLlama-13B-Instruct 6.34 5.58 4.83
CodeLlama-34B-Instruct 6.49 5.84 5.36
DeepSeek-Coder-6.7B-Instruct 6.70 6.07 6.32
WizardCoder-33B 6.85 6.26 6.75
GPT-3.5-Turbo - 7.18 7.35
GPT-4-Turbo 8.28 - -
CodeLlama-7B-Instruct 6.00 5.46 4.72
+SFT 6.51 5.83 5.84
+DPO 7.15 6.79 5.08
+SFT+DPO 7.36 7.08 5.85
Table 5. Pass@$k$ of CodeLlama-7B-Instruct variants on HumanEval and HumanEval+.
HumanEval HumanEval+
$k$=1 $k$=10 $k$=1 $k$=10
CodeLlama-7B-Instruct 37.9 60.4 33.2 54.9
+SFT 51.2 82.9 45.6 79.3
+DPO 42.3 80.5 35.8 70.1
+SFT+DPO 43.1 75.6 36.7 69.5
6. Discussion
In this paper, we have demonstrated how LLM-as-a-Judge enables the creation of new datasets and benchmarks like CodeUltraFeedback and CodeUltraFeedback-Bench to facilitate the tuning of LLMs for better alignment with coding preferences and assess their performance. This section offers further insights from our experiments, potential improvements for LLM alignment, and the significance of exploring varied LLM-as-a-Judge configurations. Additionally, we outline some limitations of CodeUltraFeedback and CodeUltraFeedback-Bench.
6.1. LLM Alignment and LLM-as-a-Judge
Balancing Learning Objectives. A crucial area requiring deeper investigation is the balance between functional correctness and non-functional requirements. In our experiments, we demonstrate that CodeLlama-7B-Instruct+SFT+DPO shows higher performance on HumanEval/HumanEval+ compared to the base model. Understanding how to integrate functional and non-functional learning objectives could significantly advance the tuning process of LLMs and further enhance their performance across a broad range of benchmarks.
CodeUltraFeedback for Critic LLM Training. In this work, we demonstrate the utility of CodeUltraFeedback for LLM preference tuning using RLAIF and DPO. One of the drawbacks of CodeUltraFeedback-Bench is the need to leverage models like GPT-4 to evaluate other LLMs, which might turn out to be cost-prohibitive. Nonetheless, we believe the AI feedback data in CodeUltraFeedback, e.g., ratings and rationales, can be leveraged to fine-tune a small critic LLM trained to evaluate other LLMs. Prior work including Shepherd (Wang et al., 2023) and Prometheus (Kim et al., 2023) have been proposed around that idea and showed promising results.
Parallel Findings with Zephyr Our findings draw parallels with Zephyr ’s (Tunstall et al., 2023) insights on DPO, with models beginning to overfit after a few epochs, achieving 100% training set accuracy. In our case, overfitting after five DPO training epochs correlated with improved performance. Although the model overfits, the gap between chosen and rejected rewards continued to grow, suggesting that monitoring reward accuracy might not be relevant for model selection. Therefore, more experimentation on DPO could illuminate better model tuning and selection approaches for improved model alignment.
Potential Judgment Biases. We highlight GPT-3.5-Turbo’s robustness in judgment in both our dataset and benchmark. The performance ranking of LLMs remained consistent, underscoring the reliability of GPT-3.5-Turbo’s evaluations. Despite this, there remains the possibility of inherent biases in LLM judgments, such as a preference for lengthy or verbose responses. The grading procedures employed for CodeUltraFeedback and CodeUltraFeedback-Bench were carefully designed, drawing on established practices to preclude such biases (Zheng et al., 2023a; Li et al., 2023b). Single-answer grading with a reference response to the instruction and chain-of-thought prompting enables both GPT-3.5-Turbo and GPT-4-Turbo to produce consistent judgements. However, further exploration into the potential biases influencing LLM judgments is warranted.
Exploring Alternative Judges. Future studies could benefit from incorporating a wider range of judges to understand how different LLMs’ evaluative perspectives might influence the judgment process and outcomes. Additionally, the reliance on closed-source models can be mitigated by tuning small and open-source critic LLMs.
6.2. Limitations
Our methodology presupposes the existence of a highly capable model, such as GPT-3.5 or GPT-4, as a proxy to humans to accurately evaluate other LLMs. The reliance on these models is predicated upon previous findings, which indicate that GPT-3.5 and GPT-4 align closely with human judgments (Zheng et al., 2023a; Li et al., 2023b), achieving agreement rates on par with those between humans themselves on MT-Bench (Zheng et al., 2023a). Additionally, recent work (Zhuo, 2024) demonstrates that ICE-Score, a novel metric for code generation based on LLM-as-a-Judge is superior to prior metrics and has a higher correlation with human judgements. Such evidence supports the utilization of LLMs as evaluators. Although our work does not include a comparison between GPT-3.5/4 and human judgements, we hypothesize their correlation on CodeUltraFeedback-Bench to be high and leave the comparison for future work.
Our current iteration of CodeUltraFeedback-Bench relies on randomly selected samples. Future versions aim to employ a more rigorous filtering process, possibly incorporating human annotations, to enhance the quality of selected samples.
Lastly, our grading procedure in CodeUltraFeedback-Bench necessitates reference responses. This usage of reference responses gives the judge LLMs a consistent point of comparison, enabling fair and consistent judgements across LLMs outputs. Future work might explore ways to reduce this dependency.
7. Related Work
We differentiate our work from the existing landscape of datasets and benchmarks in automated software engineering, with a particular focus on execution-based benchmarks, closed-solutions datasets, and benchmarks, as well as those assessing non-functional aspects of generated code.
7.1. Execution-based Benchmarks
This category of benchmarks emphasizes the functional correctness of generated code by evaluating whether it passes a set of unit tests. Execution-based benchmarks include HumanEval (Chen and others., 2021), MBPP (Austin et al., 2021), APPS (Hendrycks et al., 2021), and DS-1000 (Lai et al., 2023), which have extensively been used to compare LLMs in code generation scenarios (Roziere et al., 2023; Muennighoff et al., 2023; Luo et al., 2023; Wei et al., 2023). Subsequent work has expanded HumanEval and MBPP benchmarks to multiple programming languages (Athiwaratkun et al., 2022; Cassano et al., 2023; Zheng et al., 2023b), more unit tests (Liu et al., 2023b), and more tasks including code summarization (Athiwaratkun et al., 2022), code repair (Muennighoff et al., 2023), code explanation (Muennighoff et al., 2023), and code translation (Zheng et al., 2023b). This category also encompasses benchmarks designed around specific use cases, such as Odex (Zhou et al., 2022), which deals with Python open-domain problems and includes manual annotations of user intents, and StudentEval (Babe et al., 2023), which utilizes student-defined prompts for Python code generation. xCodeEval (Khan et al., 2023) benchmark extends the scope to code understanding, generation, and retrieval across numerous tasks. Lastly, benchmarks like ClassEval (Du et al., 2023) and CoderEval (Yu et al., 2024) shift the focus towards evaluating LLMs’ proficiency in generating functionally correct code across various programming abstractions, such as class, file, and project levels.
Our work differs from this category of benchmarks and focuses on aligning LLMs to coding preferences and measuring their alignment using an LLM as a proxy for human evaluation without relying on unit tests or automated metrics. Nonetheless, we believe evaluating the functional correctness of LLM-generated code remains a pivotal concern complementary to the aim of CodeUltraFeedback and CodeUltraFeedback-Bench.
7.2. General Datasets and Closed-Solutions Benchmarks
Initial datasets like GitHub Java Corpus (Allamanis and Sutton, 2013), Py150/Js150 (Raychev et al., 2016a, b), and CodeSearchNet (Husain et al., 2019) laid the groundwork for evaluating language models in code-related tasks, including modelling, summarization, and search. Subsequent developments introduced benchmarks like CodeXGlue (Lu et al., 2021), CodeNet (Puri et al., 2021), XLCoST (Zhu et al., 2022), ReCode (Wang et al., 2022b), and CrossCodeBench (Niu et al., 2023), each expanding the evaluation scope to include code understanding, generation, and robustness against perturbations across various coding tasks and languages. Recently, CrossCodeEval (Ding et al., 2023) broadened this scope by assessing generative capabilities and the use of cross-file information for project-wide coding tasks.
These datasets and benchmarks evaluate language models’ core capabilities, mainly through task-specific transfer learning, relying on ground truth solutions which may overlook valid code variations. In contrast, CodeUltraFeedback and CodeUltraFeedback-Bench focus on aligning LLMs with human coding preferences. Additionally, the LLM-as-a-Judge approach serves as an alternative to automated evaluations, prioritizing nuanced assessment of natural and programming languages over strict adherence to ground truth.
7.3. Non-Functional Evaluation of LLM-Generated Code
Recent advancements have expanded LLM evaluation to include non-functional requirements (Yang et al., 2024), an aspect overlooked in earlier studies. Examples include efforts to assess the quality of AI-generated code (Liu et al., 2023a; Siddiq et al., 2023) by using static analysis tools. In contrast, our work adopts a broader strategy, employing LLMs’ advanced reasoning to tune and assess their alignment with human coding preferences. Furthermore, Yetiştiren et al. (Yetiştiren et al., 2023) evaluated LLM-generated code using quality metrics such as security, reliability, and maintainability on the HumanEval benchmark. Our work encompasses a broader evaluation scope, including more complex instructions, and CodeUltraFeedback to tune and align LLMs to coding preferences. In contrast to CyberSecEval (Bhatt et al., 2023) and EffiBench (Huang et al., 2024), which concentrate on particular code aspects like security and efficiency, our methodology based on LLM-as-a-Judge provides a holistic evaluation across several dimensions, offering a framework that can be adapted to assess an array of coding preferences. NoFunEval (Singhal et al., 2024) stands as the most closely related benchmark to our work, evaluating LLM-generated code’s non-functional properties across 958 coding problems using functional specifications and static analysis tools. In contrast, we leverage LLMs’ evaluative capabilities for LLM evaluation, overcoming static analysis tools’ limitations in capturing language nuances. Additionally, we introduce a comprehensive dataset for tuning LLM preferences, distinguishing our work as a unique contribution to the field of code LLM evaluation.
8. Conclusion and Future Work
In this paper, we introduce CodeUltraFeedback and CodeUltraFeedback-Bench, a preference dataset and benchmark of instructions for LLM alignment to five coding preferences. Our analysis reveals significant alignment disparities among various LLMs compared to GPT-3.5 and GPT-4. Furthermore, we demonstrate how CodeUltraFeedback facilitates preference tuning using SFT, RLAIF and DPO. Our experiments conclude that CodeLlama-7B-Instruct tuned with SFT and DPO outperforms 34B LLMs on CodeUltraFeedback-Bench and enhances functional correctness on HumanEval+. We hope CodeUltraFeedback and CodeUltraFeedback-Bench can further support research related to LLM alignment in automated software engineering. In future work, we plan to explore more judges to assess LLMs on CodeUltraFeedback-Bench and use CodeUltraFeedback to fine-tune our own judge LLM, thereby reducing our dependence on closed-source models like GPT-4.
Appendix A Assessment Templates
Table 6. Instruction following assessment template.
Evaluate the assistant’s fidelity to provided instructions. Assess how accurately the assistant’s responses align with user directives, noting any deviations and their justification. Evaluation Criteria:
   Precision in Following Instructions: Does the assistant adhere to the specifics of the provided instructions?
   Justification for Deviations: If deviations occur, are they justified by critical necessity or explicit user request?
   Alignment with User Directives: How well do the assistant’s responses match the user’s specified needs and expectations?
   Necessity of Deviations: Are any deviations from instructions made only in situations deemed absolutely necessary or upon direct user request?
Scoring: Rate outputs on a scale of 1 to 5:
  1. Non-Compliant: The assistant frequently deviates from instructions without necessity or user consent.
  2. Minimally Compliant: The assistant shows some adherence to instructions but deviates without strong justification.
  3. Moderately Compliant: The assistant generally follows instructions, with deviations occurring but justified by necessity or user request.
  4. Highly Compliant: The assistant closely follows instructions, with few deviations, all of which are well justified.
  5. Fully Compliant: The assistant exhibits perfect adherence to instructions, with deviations only when critically necessary and explicitly approved by the user.
Table 7. Code explanation assessment template.
Evaluate the clarity and depth of explanations accompanying code segments. Assess how well the explanation helps in understanding the code’s purpose, logic, and design choices. Evaluation Criteria:
   Clarity: How easy is it to understand the explanation?
   Depth: Does the explanation cover the logic, structure, and decisions behind the code?
   Relevance: Is the explanation relevant to the code’s purpose and design philosophy?
   Accessibility: Can a broad audience understand the explanation, regardless of their technical background?
Scoring: Rate outputs on a scale of 1 to 5:
  1. Inadequate: Explanation is unclear, superficial, or missing.
  2. Basic: Explanation covers fundamental points but lacks depth or clarity.
  3. Good: Explanation is clear and somewhat detailed, but may miss some deeper insights.
  4. Very Good: Explanation is clear, detailed, and covers most logic, structure, and decisions.
  5. Excellent: Explanation is exceptionally clear, in-depth, and makes the code’s purpose, design, and logic accessible to all users.
Table 8. Code complexity and efficiency assessment template.
Evaluate the solutions and code provided by the assistant for their time efficiency and resource management. Assess how well the code optimizes computational time and resources while ensuring the accuracy and effectiveness of the implemented algorithms. Evaluation Criteria:
   Time Efficiency: Does the code minimize computational time?
   Resource Efficiency: Does the code use resources (e.g., memory, CPU) judiciously?
   Algorithm Effectiveness: Are the chosen algorithms accurate and efficient in achieving the desired outcomes?
   Optimization: Has the code been optimized for quick processing without compromising the solution’s correctness or efficiency?
Scoring: Rate outputs on a scale of 1 to 5:
  1. Inefficient: The code is resource-heavy and slow, with little evidence of optimization for time efficiency.
  2. Somewhat Efficient: The code shows some effort towards efficiency, but significant improvements are needed in time and resource management.
  3. Moderately Efficient: The code balances time and resource use reasonably well, with effective algorithm selection, but there’s room for optimization.
  4. Highly Efficient: The code demonstrates strong time and resource efficiency, with well-chosen algorithms that provide swift and accurate results.
  5. Optimally Efficient: The code exemplifies the best practices in time and resource efficiency, with optimal algorithm selection and execution that ensure maximum effectiveness with minimal resource expenditure.
Table 9. Coding style assessment template.
Evaluate the coding style of provided code segments. Assess how well the code adheres to the best practices of the language, focusing on readability, maintainability, and efficiency in line with the language’s idiomatic style. Evaluation Criteria:
   Readability: Is the code easy to read and understand?
   Maintainability: Can the code be easily modified or extended?
   Efficiency: Does the code execute tasks in an efficient manner?
   Adherence to Idiomatic Style: Does the code follow the stylistic norms and conventions specific to the programming language?
Scoring: Rate outputs on a scale of 1 to 5:
  1. Non-Adherent: The code largely ignores language conventions, resulting in poor readability and maintainability.
  2. Somewhat Adherent: The code makes some effort to follow language conventions but lacks consistency in readability, maintainability, or efficiency.
  3. Moderately Adherent: The code is generally in line with language conventions, offering fair readability and maintainability, with room for improvement in efficiency.
  4. Highly Adherent: The code strongly adheres to language conventions, demonstrating good readability, maintainability, and efficiency.
  5. Exemplary Adherent: The code exemplifies the best practices of the language, excelling in readability, maintainability, efficiency, and idiomatic style.
Table 10. System prompt for single-answer grading.
Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user instruction displayed below. You will be given a Reference Response and the Assistant’s Response. Begin your evaluation by comparing the Assistant’s Response with the Reference Response. Your overall evaluation needs to be reflective of the specified Evaluation Criteria. Be as objective as possible. After providing your rationale, you must rate the Assistant’s Response on a scale of 1 to 10. DO NOT give a response to the instruction, ONLY provide your rationale followed by the rating.
Table 11. Instruction following template for single-answer grading.
# Instruction Following Assessment
Evaluation Criteria:
  - Precision in Following Instructions: Does the assistant adhere to the specifics of the provided instructions?
  - Justification for Deviations: If deviations occur, are they justified by critical necessity or explicit user request?
  - Alignment with User Directives: How well do the assistant’s responses match the user’s specified needs and expectations?
  - Necessity of Deviations: Are any deviations from instructions made only in situations deemed absolutely necessary or upon direct user request?
## Format:
### Instruction
  [Instruction]
### Reference Response:
  [Reference]
### Assistant’s Response:
  [Response]
### Output
Rationale: [Rationale for the rating in short sentences]
Rating: [Rating for Assistant’s Response]
## Example of output:
Rationale: The assistant’s response …
Rating: 2/10
## Annotation
### Instruction
  {instruction}
### Reference Response:
  {reference}
### Assistant’s Response:
  {response}
### Output
Table 12. Code explanation template for single-answer grading.
# Code Explanation Quality Assessment
Evaluation Criteria:
  - Clarity: How easy is it to understand the explanation?
  - Depth: Does the explanation cover the logic, structure, and decisions behind the code?
  - Relevance: Is the explanation relevant to the code’s purpose and design philosophy?
  - Accessibility: Can a broad audience understand the explanation, regardless of their technical background?
## Format:
### Instruction
  [Instruction]
### Reference Response:
  [Reference]
### Assistant’s Response:
  [Response]
### Output
Rationale: [Rationale for the rating in short sentences]
Rating: [Rating for Assistant’s Response]
## Example of output:
Rationale: The assistant’s response …
Rating: 2/10
## Annotation
### Instruction
  {instruction}
### Reference Response:
  {reference}
### Assistant’s Response:
  {response}
### Output
Table 13. Code complexity and efficiency template for single-answer grading.
# Code Complexity and Efficiency Assessment
Evaluation Criteria:
  - Time Efficiency: Does the code minimize computational time?
  - Resource Efficiency: Does the code use resources (e.g., memory, CPU) judiciously?
  - Algorithm Effectiveness: Are the chosen algorithms accurate and efficient in achieving the desired outcomes?
  - Optimization: Has the code been optimized for quick processing without compromising the solution’s correctness or efficiency?
## Format:
### Instruction
  [Instruction]
### Reference Response:
  [Reference]
### Assistant’s Response:
  [Response]
### Output
Rationale: [Rationale for the rating in short sentences]
Rating: [Rating for Assistant’s Response]
## Example of output:
Rationale: The assistant’s response …
Rating: 2/10
## Annotation
### Instruction
  {instruction}
### Reference Response:
  {reference}
### Assistant’s Response:
  {response}
### Output
Table 14. Code readability template for single-answer grading.
# Code Readability Assessment
Evaluation Criteria:
  - Clarity: How clear and understandable are the code and its accompanying comments/documentation?
  - Conciseness: Are the comments and documentation succinct yet informative?
  - Relevance: Do the comments and documentation directly contribute to explaining the code’s logic, objectives, and functionality?
  - Comprehensibility: Can users of varying technical backgrounds easily grasp the code’s purpose and how it works?
## Format:
### Instruction
  [Instruction]
### Reference Response:
  [Reference]
### Assistant’s Response:
  [Response]
### Output
Rationale: [Rationale for the rating in short sentences]
Rating: [Rating for Assistant’s Response]
## Example of output:
Rationale: The assistant’s response …
Rating: 2/10
## Annotation
### Instruction
  {instruction}
### Reference Response:
  {reference}
### Assistant’s Response:
  {response}
### Output
Table 15. Coding style template for single-answer grading.
# Coding Style Assessment
Evaluation Criteria:
  - Readability: Is the code easy to read and understand?
  - Maintainability: Can the code be easily modified or extended?
  - Efficiency: Does the code execute tasks in an efficient manner?
  - Adherence to Idiomatic Style: Does the code follow the stylistic norms and conventions specific to the programming language?
## Format:
### Instruction
  [Instruction]
### Reference Response:
  [Reference]
### Assistant’s Response:
  [Response]
### Output
Rationale: [Rationale for the rating in short sentences]
Rating: [Rating for Assistant’s Response]
## Example of output:
Rationale: The assistant’s response …
Rating: 2/10
## Annotation
### Instruction
  {instruction}
### Reference Response:
  {reference}
### Assistant’s Response:
  {response}
### Output
Table 16. Instruction following principles.
The assistant is required to strictly follow the provided instructions, making sure that all responses and actions are in exact accordance with user directives. Only in critical situations or upon direct user request should it deviate from these instructions.
It’s essential for the assistant to rigorously stick to the instructions given, aligning its responses and actions accurately with user demands. Deviations are permissible solely when indispensable or specifically asked for by the user.
The assistant must ensure precise compliance with instructions, aligning its responses and actions directly with user requirements. Any deviation should occur only if absolutely necessary or if explicitly requested by the user.
In following instructions, the assistant should adhere closely to what is asked, ensuring responses and actions match the user’s specified needs. Departure from these instructions is allowed only when critically needed or when the user explicitly allows it.
The assistant is obliged to faithfully follow the instructions provided, ensuring its responses and actions mirror user expectations precisely. Only in necessary circumstances or at the user’s express request should it deviate.
The assistant needs to maintain strict adherence to user instructions, aligning its responses and actions specifically with the user’s requests. It should deviate from these instructions only when there is an absolute necessity or upon user’s explicit instruction.
It is crucial for the assistant to be diligent in following user instructions, keeping its responses and actions in tight alignment with what is requested. Deviations should be considered only under essential conditions or if requested by the user.
The assistant’s responses and actions must conform closely to the instructions given by the user. It should only diverge from these instructions in situations of utmost necessity or when explicitly asked by the user.
The assistant should execute tasks in strict accordance with user instructions, ensuring all actions and responses are directly in line with user requirements. Only in pressing situations or with clear user consent should it deviate.
For the assistant, precise adherence to the given instructions is mandatory, with responses and actions reflecting user requests accurately. Deviating from these guidelines is acceptable only in critical instances or when the user specifically requests it.
Table 17. Code explanation principles.
The assistant is expected to provide detailed and accessible explanations for its code, clarifying the reasoning, architecture, and decisions involved, thereby aiding users in comprehending both the function and the rationale behind the code.
It’s important for the assistant to deliver comprehensive and intelligible explanations for the code, breaking down the logic, framework, and choices, to help users grasp the code’s purpose and construction.
The assistant must ensure that its code comes with clear, in-depth explanations, illuminating the underlying logic, structure, and choices, enabling users to not only understand the code’s workings but also its design philosophy.
In providing code, the assistant should also offer exhaustive and lucid explanations, elucidating the thought process, organization, and decisions behind the code, thus facilitating user understanding and application.
The assistant should accompany its code with explicit and detailed explanations, shedding light on the logic, configuration, and decisions, helping users to comprehend and utilize the code effectively.
For every piece of code, the assistant must provide clear, comprehensive explanations, detailing the logic, structure, and reasoning, enabling users to fully understand and apply the code in their contexts.
The assistant is obliged to offer clear-cut and extensive explanations with its code, dissecting the logic, structure, and strategic choices, ensuring users gain a complete understanding of the code’s purpose and design.
Alongside its code, the assistant should present thorough and straightforward explanations, clarifying the underlying logic, framework, and choices, aiding users in understanding and adapting the code efficiently.
The assistant must ensure that each code segment is accompanied by transparent and thorough explanations, unraveling the logic, structure, and choices, thus empowering users to grasp and leverage the code for their needs.
It is essential for the assistant to couple its code with lucid, detailed explanations, unfolding the logic, composition, and decisions within the code, to help users not only understand its functionality but also the rationale behind its creation.
Table 18. Code complexity and efficiency principles.
The assistant is tasked with delivering solutions and code that prioritize time efficiency, focusing on reducing computational duration and resource expenditure while ensuring the accuracy and effectiveness of the algorithms used.
It’s essential for the assistant to offer solutions and code that are optimized for quick processing, designed to minimize the time and resources required for execution, while maintaining correctness and efficiency.
The assistant must ensure that its solutions and code are developed with optimal time efficiency in mind, aiming to lessen computational time and resource use, and selecting algorithms that are both accurate and swift in execution.
In providing solutions, the assistant should focus on achieving the highest time efficiency, minimizing the amount of computational time and resources needed, while ensuring the selected algorithms are both correct and efficient.
The assistant’s goal should be to present solutions and code that excel in time efficiency, designed to curtail the duration of computation and resource usage, while adhering to algorithms that are effective and precise.
For every solution and piece of code, the assistant is expected to prioritize time efficiency, striving to reduce computational time and resources, and favoring methods that are not only accurate but also swift.
The assistant needs to concentrate on offering solutions and code that are time-efficient, minimizing computational and resource demands, while choosing algorithms that are both effective and efficient.
In its solutions and code, the assistant should aim for optimal time efficiency, designing methods that cut down on computational time and resource use, while ensuring the approaches used are both correct and swift.
The assistant is required to prioritize time efficiency in its solutions and code, focusing on minimizing the time and resources involved in computation, and selecting algorithms that are not only accurate but also quick.
The assistant’s objective should be to provide solutions and code that excel in time efficiency, aiming to reduce both computational time and resource consumption, while utilizing algorithms that are effective and efficient in their operation.
Table 19. Code readability principles.
The assistant must ensure its code and responses are marked by ease of understanding, incorporating straightforward, brief comments and documentation that clarify the code’s logic, purpose, and operational process, facilitating user comprehension and learning.
It is crucial for the assistant to make its code and responses readily comprehensible, using clear and concise explanations in comments and documentation to illuminate the reasoning, objectives, and functions of the code, thus aiding in user education and application.
The assistant should provide code and responses that are easily graspable, with lucid, succinct comments and documentation elucidating the underlying logic, goals, and functionality of the code, enhancing the user’s ability to follow and absorb the material.
In delivering code and responses, the assistant is expected to prioritize understandability, featuring clear, compact comments and documentation that detail the code’s logic, intent, and capabilities, thereby assisting users in their learning journey.
The assistant is tasked with ensuring that both its code and responses are straightforward to comprehend, integrating clear, brief comments and documentation that explicate the code’s reasoning, purpose, and mechanisms, thus streamlining the learning process for users.
For effective readability, the assistant should ensure its code and responses are effortlessly intelligible, employing clear, concise comments and documentation that convey the logic, objectives, and functionality of the code, simplifying the learning curve for users.
The assistant’s code and responses need to be accessible in terms of understanding, necessitating the use of clear, brief comments and documentation that outline the logic, intent, and function of the code, thereby enhancing user comprehension and learning.
The assistant has the responsibility to ensure its code and responses are straightforward to understand, necessitating lucid, concise comments and documentation that outline the code’s reasoning, purpose, and operational aspects, key to ensuring users can easily follow and benefit from the provided material.
The assistant should maintain a high level of readability in both its code and responses, achieved by integrating clear, straightforward comments and documentation that illuminate the code’s reasoning, purpose, and functionality, to assist users in easily grasping and learning from the content.
In providing code and responses, the assistant must focus on making them easily understandable, entailing the inclusion of lucid, succinct comments and documentation that explicate the code’s logic, aim, and functional aspects, essential for ensuring users can easily follow and benefit from the provided content.
Table 20. Coding style principles.
The assistant should write code that exemplifies the best practices of the language used, focusing on readability, maintainability, and efficiency in line with the language’s idiomatic style.
It is important for the assistant to adopt a coding style that is characteristic of the language, balancing clarity, conciseness, and efficiency to produce easily understandable and efficient code.
The assistant’s coding approach should reflect the standard conventions of the language, prioritizing a style that is both effective and in harmony with the language’s best practices.
In coding, the assistant must aim for a style that is not only syntactically correct but also adheres to the idiomatic nuances of the programming language, ensuring efficiency and readability.
The assistant is expected to craft code that follows the established stylistic norms of the programming language, optimizing for a balance between clarity, efficiency, and language-specific best practices.
The assistant should maintain a coding style that aligns with the language’s conventional practices, ensuring that the code is efficient, readable, and easily maintained.
Coding in a style that is representative of the language’s community standards, the assistant should strive for code that is succinct, efficient, and easily interpreted by others.
The assistant’s coding should demonstrate a clear adherence to the stylistic principles of the language, focusing on creating code that is both high-performing and elegantly written.
In its coding practices, the assistant should embody the ideal balance of efficiency, readability, and adherence to the language’s stylistic conventions, ensuring optimal code performance.
The assistant’s coding style should be a reflection of the language’s standard practices, emphasizing a clean, efficient, and idiomatic approach to ensure both performance and readability.